OpenStudy (anonymous):

What is the solution to the equation? 7x + 6 = –22 A. B. x = –4 C. D. x = 4

OpenStudy (anonymous):

@dtan5457 @confluxepic @CrazyCountryGirl

OpenStudy (dtan5457):

b. x=-4 subtract the 6 and you get 7x=-28 x=-4

OpenStudy (anonymous):

What is the solution to the equation? 10y – 25 = 5 A. y = –3 B. y = –2 C. y = 2 D. y = 3

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

y = 3 add the 25 to the other sides 10y=30 y=3

OpenStudy (anonymous):

What is the solution to the equation? 05_05_LQ_3.gif A. z = 50 B. z = 44 C. z = –2 D. z = – 40

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

the gif doesnt work

OpenStudy (anonymous):

re fresh the page

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

it's unclickable better yet why is it a gif

OpenStudy (anonymous):

their you should be able to see it their

OpenStudy (dtan5457):

it's -40

OpenStudy (dtan5457):

-40/5=-8 double negatives make it positive 1+8=9

OpenStudy (anonymous):

What is the solution to the equation? 6 + 2s – 8s = 18 A. s = –4 B. s = –2 C. s = 2 D. s = 4

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

s = –2

OpenStudy (dtan5457):

plug it in and you'll see it equals 18

OpenStudy (anonymous):

What is the solution to the equation? –4q + 8 = –2q – 16 A. q = –12 B. q = –4 C. q = 4 D. q = 12

OpenStudy (anonymous):

last one

OpenStudy (dtan5457):

q = 12 first term is -48+8=-40 2nd term is -24-16=-40

OpenStudy (anonymous):

thank you so much!!